dakotawalker:

Melissa and Emily are playing at the pool. They have three different measuring jars for liters, cups, and pints. Melissa poured 7 cups of water and 3 liters of water into the pint jar and it was filled up to 9.8 pints. Later, Emily started with 5 liters of water in the pint jar and took out 9 cups. The remaining water level was equal to 6 pints. Model the given situation as a system of linear equations and solve it for liters. Based on your solution, what do you notice about the relationship between pints and cups, and pints and liters?

Hero:

Hi @dakotawalker welcome to QC. Were you able to start this one? If so, how far have you gotten with solving the system?

dakotawalker:

hi hero. i have tried to start it but i don't really understand it.

Hero:

Just a moment. Working on it.

Hero:

Apparently the system is supposed to be \(\begin{align*} 7C + 3L &= 9.8P\\ -9C + 5L &= 6P \end{align*}\)

Hero:

In order to solve this, we have to eliminate one of the variables.

Hero:

We have to solve for L so we can't eliminate that variable. We have to eliminate either C or P.

Hero:

Are you still with me @dakotawalker

dakotawalker:

yes please keep going

Hero:

So I think it's easier to eliminate C. How do you suggest that we go about doing this?

dakotawalker:

would combining them then subtracting work

Hero:

Well, no, that's not how elimination works. We need to do something to each equation so that the coefficients of the C terms have opposite values. That way we can add the equations together to eliminate C

Hero:

Any idea of what we can do to accomplish that?

dakotawalker:

we need to make the c terms have the same numbers but be either positive or negative

Hero:

And how would we go about doing that?

dakotawalker:

we do seven times nine for both leaving one as a negative

Hero:

The correct way to describe the process is to say: We multiply both sides of one equation by 7 and multiply both sides of the other equation by 9

Hero:

And that process looks like this: \(\begin{align*} 9(7C + 3L) &= 9(9.8P)\\ 7(-9C + 5L) &= 7(6P) \end{align*}\)

Hero:

Would you mind doing the next step?

dakotawalker:

do we divide or do make an equation

Hero:

Are you familiar with the distributive property? a(b + c) = ab + ac. That's what we apply here. Also if you see a(b) it means multiply a and b.

dakotawalker:

so it will be 9(9.8P+6P)= 9(9.8P)+9(6P)

Hero:

Can you show me how you're applying the distributive property to arrive at this result?

Hero:

I meant for you to apply the distributive property to this system: \(\begin{align*} 9(7C + 3L) &= 9(9.8P)\\ 7(-9C + 5L) &= 7(6P) \end{align*}\)

Hero:

And we only apply it on the LHS. Are you familiar with the distributive property?

dakotawalker:

A little but i have trouble with it and it has been a long time since i used it

Hero:

I figured. You should review distributive property before continuing. I recommend YouTube and Khan Academy for reviewing and drilling on this concept.

dakotawalker:

ok thank you but i have to go